OpenStudy (anonymous):

please help definite integral from 1 to 2 lnx/x^2dx

OpenStudy (turingtest):

integration by parts I believe...

OpenStudy (turingtest):

\[u=\ln x\]\[dv=x^{-2}\]

OpenStudy (dumbcow):

yep that will work du = 1/x v = -1/x \[\rightarrow \int\limits_{?}^{?}\frac {\ln x}{x^{2}}dx = -\frac{\ln x}{x} +\int\limits_{?}^{?}\frac{1}{x^{2}} dx\]